OpenStudy (anonymous):

a certain game consists of rolling a single fair die and pays off as follows: $5 for a 6, $2 for a 5, $1 for a 4, and no payoff otherwise. find the expected winnings for this game.

jimthompson5910 (jim_thompson5910):

I'm assuming you don't pay to play this game right?

OpenStudy (anonymous):

exactly

jimthompson5910 (jim_thompson5910):

if so, then P(rolling 6) = 1/6 W(rolling 6) = 5 Note: W(rolling 6) = "amount of money you win for rolling a six...so it's your winnings if you roll a 6" P(rolling 5) = 1/6 W(rolling 5) = 2 P(rolling 4) = 1/6 W(rolling 4) = 1 P(rolling 3,2,or1) = 3/6 W(rolling 3,2,or1) = 0

jimthompson5910 (jim_thompson5910):

Expected Value E[X] = P(rolling 6)*W(rolling 6)+P(rolling 5)*W(rolling 5)+P(rolling 4)*W(rolling 4)+P(rolling 3,2,1)*W(rolling 3,2,1) E[X] = (1/6)*(5)+(1/6)*(2)+(1/6)*(1)+(3/6)*(0) E[X] = 5/6 + 2/6 + 1/6 + 0/6 E[X] = (5 + 2 + 1 + 0)/6 E[X] = 8/6 E[X] = 4/3 E[X] = 1.33333333... E[X] = 1.33

jimthompson5910 (jim_thompson5910):

So assuming you do not pay to play, you would expect to earn $1.33 for each roll (on average)